OpenStudy (anonymous):

find the distance between p1 (3,-195 degrees) and p2 (-4,-94 degrees) on the polar plane. Round your answer to the nearest thousandth.

OpenStudy (michele_laino):

here we can use the standard formula, namely: \[ \large d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} \]

OpenStudy (anonymous):

x=1/8 @Michele_Laino

OpenStudy (michele_laino):

we have to use the subsequent formulas: \[\left\{ \begin{gathered} x = r\cos \theta \hfill \\ y = r\sin \theta \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

for example I can write this: \[\begin{gathered} {x_1} = 3\cos ( - 195) = 3\cos (195) = ...? \hfill \\ {y_1} = 3\sin ( - 195) = - 3\sin (195) = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

similarly for the second point: \[\begin{gathered} {x_2} = 4\cos ( - 94) = 4\cos (94) = ...? \hfill \\ {y_2} = 4\sin ( - 94) = - 4\sin (94) = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

is it 4.519

OpenStudy (michele_laino):

I got 5.245